OpenStudy (anonymous):

Series Solution Help y" + sin(x)y = 0

OpenStudy (anonymous):

Find the fundamental set of solutions for the following equation\[y"+(\sin(x))y = 0\]

OpenStudy (anonymous):

Actually, just the first four nonzero terms.

OpenStudy (therealmeeeee):

i will write down the list ov zeros

OpenStudy (therealmeeeee):

Assume that y = a₀ + a₁x + a₂x^2 + a₃x^3 + a₄x^4 + ... ==> y'' = 2a₂ + 6a₃x + 12a₄x^2 + ... Substituting this into the DE (along with the series for sin x) yields [2a₂ + 6a₃x + 12a₄x^2 + ...] + [x - x^3/3! + ...] [a₀ + a₁x + a₂x^2 + a₃x^3 + a₄x^4 + ...] = 0 ==> [2a₂ + 6a₃x + 12a₄x^2 + ...] + [a₀x + a₁ x^2 + (-a₀/3! + a₂) x^3 + ...] = 0, via multiplication ==> 2a₂ + (a₀ + 6a₃) x + (a₁ + 12a₄) x^2 + ... = 0. Equating like coefficients (assuming that a₀ and a₁ are arbitrary): 2a₂ = 0 ==> a₂ = 0 a₀ + 6a₃ = 0 ==> a₃ = -a₀/6 a₁ + 12a₄ = 0 ==> a₄ = -a₁/12. Hence, y = a₀ + a₁x + 0x^2 + (-a₀/6)x^3 + (-a₁/12)x^4 + ... ..............= a₀ [1 - (1/6)x^3 + ...] + a₁ [x - (1/12)x^4 + ...]. I hope this helps!

OpenStudy (therealmeeeee):

well do it help??????????????????????????

OpenStudy (therealmeeeee):

do it helppppppp

OpenStudy (anonymous):

Wonderful. Yes, then I am deriving it correctly. Thank you for your response. So, every time we have a product of summations, we can just multiply term by term?

OpenStudy (therealmeeeee):

yes u can

OpenStudy (therealmeeeee):

Click Best Response

OpenStudy (mathmale):

I take it that you're studying differential equations, Mr. Bauer (or is it Ms. Bauer?). I'd classify y'' = y*sin x as a SEPARABLE differential equation. Real Mee has come up with an elegant approach involving series. As an alternative approach: I'd rewrite y'' as (dy/dx); thus, (dy/dx) = y*sin x. Separating variables, (dy/y) = (sin x)dx. Integrating, ln |y| = -cos x + c or -cos x + ln C Evaluating e^ln|y| = e^(-cos x + ln C), we'd get |y| = (e^-cos x) times some other constant. Thus, the general solution of the d. e. would be y = e^(-cos x) times some constant.

OpenStudy (anonymous):

So, we write y" as (dy/dx) and not (d^2y/dx^2)?

OpenStudy (anonymous):

What do you think, @OOOPS ?

OpenStudy (anonymous):

I am with @TheRealMeeeee

OpenStudy (anonymous):

but not completely satisfy with the answer. However, I don't know how to solve it yet. I need look up my note

OpenStudy (anonymous):

if the coefficient of y is a polynomial, we expand that way, yours is a trig, cannot do the same

OpenStudy (anonymous):

Yes. There is some initial conditions involved and you also have to carry out a couple more terms to find out the relations of a_3 and a_4.

OpenStudy (anonymous):

Actually, never mind. We are fine as far as that goes.

OpenStudy (anonymous):

oh yea, @TheRealMeeeee did it for you, he wrote the Maclaurin series in polynomial form, my bad, I carelessly read it

OpenStudy (anonymous):

@mathmale please clarify your reasoning.

OpenStudy (mathmale):

I must apologize for having written the differential equation incorrectly. You are indeed correct: y'' can be re-written as (d^2 y / dx^2). Let me know whether you want to continue this discussion. I believe those who tried to help you find a series solution to the d. e. were more on target, but my more direct approach would at least give you a way to check any answer derived from a series solution.

OpenStudy (anonymous):

Yes, we can plug our solution back in to test our answer. So far, it is my understanding that separable equations were only for first order differential equation. It's interesting to consider whether or not there are separable equations for higher order D.Es as well.